# The specification approach to Equilibrium States for Parabolic Rational Maps

### v1

@@ 5. Positive entropy and the hyperbolicity criterion
We now apply a method by Call and the second-named author [10] to show that our equilibrium states have positive entropy. These techniques were developed further by Call and the following result is a version of [9, Theorem B]. The non-invertible case follows the same proof, with some simpler bookkeeping. The statement we give here is also simpler because we assume a global positive expansivity condition.
@@
The method of proof is simply that the criteria guarantee a unique equilibrium state for $(X\times X,f\times f,\Phi)$. Ledrappier showed that a unique equilibrium state for $(X\times X,f\times f,\Phi)$ guarantees a unique equilibrium state with the K-property for the system $(X,f,\varphi)$, see [9, Theorem 3.3]. To show that pressure gap holds for the product system, we will apply the following result by Call.
@@ Acknowledgements
We thank Roland Roeder and Liz Vivas for helpful discussions.

### v2

@@ 5. Positive entropy and the hyperbolicity criterion
We now apply a method by Call and the second-named author [15] to show that our equilibrium states have positive entropy. These techniques were developed further by Call and the following result is a version of [13, Theorem B]. The non-invertible case follows the same proof, with some simpler bookkeeping. The statement we give here is also simpler because we assume a global positive expansivity condition.
@@
The method of proof is simply that the criteria guarantee a unique equilibrium state for $(X\times X,f\times f,\Phi)$. Ledrappier showed that a unique equilibrium state for $(X\times X,f\times f,\Phi)$ guarantees a unique equilibrium state with the $K$-property (i.e. every non-trivial factor of the measure has positive entropy) for the system $(X,f,\varphi)$, see [13, Theorem 3.3]. To show that pressure gap holds for the product system, we will apply the following result by Call.
@@ Acknowledgements
We thank Roland Roeder, Liz Vivas and Juan Rivera-Letelier for helpful discussions and comments.